OpenStudy (anonymous):

Evaluate f(x) = x 2 + 1 for f(-1). 0 undefined 2

OpenStudy (anonymous):

f(x) = 2x+1 Plug in the values accordingly 2*(-1) = -2 +1 = -1. Your answer is -1.

OpenStudy (anonymous):

@iLikeCookies_ be careful?

OpenStudy (anonymous):

What..? Oh yeah, the function is -1.

OpenStudy (anonymous):

@iLikeCookies_ you are wrong.

OpenStudy (anonymous):

The function is -1, not x.

OpenStudy (anonymous):

the answer is not -1.